find the ordinate of the centroid of the area bounded by the parabola y=x^2 and the line y=2x+3
1
Expert's answer
2021-08-24T08:08:17-0400

f(x)= x2 , g(x)=2x + 3


A= "\int_a^b" (g(x) - f(x))dx


"\overline{x}" = "\frac{1}{A}" "\int_a^b" x(g(x) - f(x))dx


"\overline{y}" ="\frac{1}{A}" "\int_a^b" "\frac{1}{2}" (g(x)2 - f(x)2)dx


Find a and b;

x2= 2x+3

x2 -2x-3=0

x1=-1, x2=3


A = "\int_a^b" (2x-3-x2)dx = x2-3x-"\frac{x^3}{3}" "\vert_{-1}^3"


= 9+9−9−(1−3−"\frac{1}{3}" )=11-"\frac{1}{3}" ="\frac{34}{3}"


"\overline{x}" = "\frac{3}{34}" "\int_{-1}^3" x(2x +3 -x2)dx = "\frac{3}{34}" "\int_{-1}^3" (2x2+3x-x3)dx =


"\frac{3}{34}" ("\frac{2x^3}{3}" +"\frac{3x^2}{2}" -"\frac{x^4}{4}""\vert_{-1}^{3}" ) = "\frac{3}{34}" (18+"\frac{27}{2}"- "\frac{81}{4}" -("-\frac{2}{3} + \frac{3}{2} -\frac{1}{4}" ) ) ="\frac{16}{17}"



"\overline{y}" = "\frac{3}{34}" "\int_{-1}^3" "\frac{1}{2}" ((2x +3)2 -x4)dx = "\frac{3}{68}" "\int_{-1}^3" (4x2+12x +9-x4)dx =


"\frac{3}{68}" ("\frac{4x^3}{3}" -6x2 -9x-"\frac{x^5}{5}""\vert_{-1}^{3}" ) = "\frac{3}{68}" (36 + 54 + 27"-\frac{243}{5}-" ("-\frac{4}{3} +6-9 -\frac{1}{5}" ) ) ="\frac{16}{5}"



The coordinate of centroid is ("\frac{16}{17} , \frac{16}{5}" )